OpenStudy (anonymous):

Lily was supposed to pay Andy $6700 6 months ago and $2140 in 5 months. Instead, she wants to repay these amounts with a payment of $3100 today and the balance in 3 months. Calculate the balance if the simple interest rate is 5.7% p.a. and the agreed focal date is 3 months from now.

OpenStudy (tkhunny):

Well, we need all the values to add up in 3 months. 6700 needs to move forward 9 months. 2140 needs to move back 2 months. <== How do you do THAT?! 3100 needs to move forward 3 months. The final payment is in just the right place already.

OpenStudy (anonymous):

is this right? $6700(1+0.057x9/12)+$2140(1+0.057x2/12)^-1=$3100(1+0.057x3/12)+x x=6986.425+2119.86-3144.175 x=5962.11

OpenStudy (dumbcow):

yes i believe you are correct :)

OpenStudy (tkhunny):

Seriously, why did you doubt? Good work.

OpenStudy (anonymous):

Thank you. I'm doubting myself cause I wasn't sure if I am right or wrong.Cause teacher didn't really teach us about it.

OpenStudy (tkhunny):

Maybe a little more confidence today, than yesterday? :-)